The following data represent the prices (in rand) of seven vehicles at a dealer:

210 125; 188 623; 375 016; 214 700; 132 210; 235 910; 198 410

The mean price of these vehicles is R222 142. Now leave out the most expensive vehicle (with a price of

R375 016) and calculate the mean price for the remaining vehicles. The amount by which the mean price

per vehicle drops after the most expensive vehicle was removed, rounded to the nearest rand, is


1
Expert's answer
2021-06-10T07:25:38-0400
"\bar{x}=\dfrac{1}{6}(210 125+188 623+214 700+132 210"

"+235 910+198 410)=196663"

"222142-196663=25479"



The amount by which the mean price per vehicle drops after the most expensive vehicle was removed, rounded to the nearest rand, is "Rs\ 25479."